OpenStudy (anonymous):

An upward force acts on a proton as it moves with a speed of 2.0 × 105 meters/second through a magnetic field of 8.5 × 10-2 tesla running from east to west. What is the force and direction of the force on the proton?

OpenStudy (anonymous):

F = q v × B q= charge of proton=1.602*10^-19 Coulombs v= 2*10^5 m/s B= 8.5*10^-2 F=2.723*10^.15 N